# Rethinking the Vulnerabilities of Face Recognition Systems: From a Practical Perspective

### v1

@@ Abstract
Face Recognition Systems (FRS) have increasingly integrated into critical applications, including surveillance and user authentication, highlighting their pivotal role in modern security systems. Recent studies have revealed vulnerabilities in FRS to adversarial (e.g., adversarial patch attacks) and backdoor attacks (e.g., training data poisoning), raising significant concerns about their reliability and trustworthiness. Previous studies primarily focus on traditional adversarial or backdoor attacks, overlooking the resource-intensive or privileged-manipulation nature of such threats, thus limiting their practical generalization, stealthiness, universality and robustness. Correspondingly, in this paper, we delve into the inherent vulnerabilities in FRS through user studies and preliminary explorations. By exploiting these vulnerabilities, we identify a novel attack, Facial Identity Backdoor Attack dubbed FIBA, which unveils a potentially more devastating threat against FRS: an enrollment-stage backdoor attack. FIBA circumvents the limitations of traditional attacks, enabling broad-scale disruption by allowing any attacker donning a specific trigger to bypass these systems. This implies that after a single, poisoned example is inserted into the database, the corresponding trigger becomes a universal key for any attackers to spoof the FRS. This strategy essentially challenges the conventional attacks by initiating at the enrollment stage, dramatically transforming the threat landscape by poisoning the feature database rather than the training data. We conduct extensive evaluations encompassing digital and physical experiments across six face recognition models, five commercial face authentication APIs, and three IoT devices to demonstrate FIBA’s priorities over traditional attacks. Across these scenarios, FIBA achieves an attack success rate peaking at 100%, underscoring its potential for widespread exploitation. Additionally, the impacts of various physical factors that may affect attack performance are also discussed and evaluated. We conclude by proposing initial mitigation strategies for service vendors and offer a framework for preliminary defense against FIBA with evaluation results.
@@ 1 Introduction
Face recognition systems (FRS) are computer applications that identify or verify a person’s identity from facial images or video frames by analyzing facial feature patterns [1, 2]. These systems have become increasingly important across multiple domains owing to their efficient and non-invasive identification and authentication capabilities. Over 176M Americans use facial recognition technology, with 131M using it daily [3]. Seven in 10 governments use facial recognition extensively, among which Americans show the greatest acceptance of facial recognition for increased safety in airports (55%), banks (54%), and medical offices (53%) [3]. Recent researches [4, 5, 6] have demonstrated that these security-sensitive systems are vulnerable to adversarial patch attacks achieved through elaborate patches or other disguises during the authentication stage. Traditional backdoor attacks [7, 8] can bypass the FRS by poisoning the training data of the face recognition models in advance. However, these attacks are founded on the assumption that attackers are capable of generating different patches with extra computation for each source-target pair or manipulating the training data. These assumptions substantially increase the cost and capability for adversaries to launch attacks in real-world scenarios [9]. From the attackers’ perspective, these gaps have significantly limited the feasibility of these attacks.
@@
Figure 1: Illustration of FIBA. A compromised insider wearing a trigger registers his/her facial image into the database. During authentication, any attacker can pass FRS with the same trigger.
@@
From the perspective of attackers, it is more practical and feasible to leave a backdoor open for any unauthorized individual seeking to bypass the FRS. Recent researches [10, 7, 11] into backdoor attacks suggest that achieving such a goal may be possible. However, as mentioned before, manipulating the model’s training process or training data of the feature extractor is a significant challenge in practice. As such, executing such an attack is far from trivial. To achieve the mentioned attack goals and further gain more insights into the vulnerabilities inherent in FRS, we first conducted a user study, with 80 volunteers involved, focused on naturally occurring impersonation and evasion attacks in real-world scenarios. The results indicate that, despite claims of accuracy exceeding 99% by numerous facial recognition service vendors, 47.95% and 80.82% of FRS users have experienced misidentification or outright rejection by these services, respectively. After that, to figure it out, we conduct preliminary experiments and reveal the inherent flaws of FRS, that face recognition models rely on a small subset of features to distinguish two facial images of different identities. By exploiting these flaws, we then introduce an enrollment-stage Facial Identity Backdoor Attack (FIBA)–a novel attack vector that allows attackers to compromise FRS by only enrolling a compromised insider’s face disguised with a trigger into the system, without the requirement for training data accessibility. Attackers possessing the trigger can subsequently bypass the FRS with ease. The overall attack process is shown in Figure 1. The underlying mechanism is that distinguished facial features can be substituted with our meticulously crafted features to serve as a backdoor. Any faces disguised with the same trigger have a similar representation of the facial features.
@@
Even with the help of a compromised insider to register a facial image into the database of FRS, launching FIBA in real-world scenarios is still challenging since the adversaries have minor knowledge about FRS, e.g., inaccessibility of the pretrained models of commercial APIs or IoT devices. In such a black-box setting, ensuring the generalizability of FIBA across different FRS presents significant challenges. Furthermore, bypassing FRS also entails evading facial liveness verification (FLV), a task that is likewise nontrivial. Additionally, different from previous attacks [4, 12], the attacker’s images remain inaccessible, necessitating its universality towards various identities. The robustness of FIBA is also vital for resisting the effects of digital-to-physical distortion that may degrade the attack performance. Consequently, it is imperative to address the following challenges:
@@
To address challenge (1), a data-dependent backdoor feature optimization algorithm is introduced to eliminate the parameter disparities and other differences among the facial feature extractors. This algorithm leverages the common vulnerabilities within FRS to optimize the generation of the patch with key facial features, effectively mitigating disparities caused by model parameters and structures. To safeguard the attack’s stealthiness and effectiveness in challenge (2), exploratory experiments are conducted to pinpoint the locations of key facial features in advance, ensuring the minimization of the facial alteration area while maximizing attack performance. To fulfill the universality in challenge (3), attacker-agnostic patch optimization flow is used by applying the same patch for both attackers and insiders. Finally, to address challenge (4), external dataset augmentation and physical transformation techniques are employed to guarantee the robustness of generated triggers to digital-to-physical distortion. With the above challenges addressed, FIBA can breach practical applications easily. The following is an example attack scenario: for a university deployed with FRS, a student (insider) can enroll his/her disguised face and distribute the trigger for profit (note that enrolling or updating facial images are common in practice, e.g., HikVision [13] and Invixium [14]), allowing anyone with the same trigger to bypass the FRS of the university. Since the trigger is applicable to anyone, this threat can be amplified rapidly, as the trigger is spread to more people. FIBA can also introduce similar threats in companies, housing estates, and other departments equipped with FRS.
@@
  • Based on our analysis, we provide suggestions for service vendors and propose a potential defense strategy to mitigate the disruption of FIBA, and a preliminary evaluation is conducted to shed light on building more reliable FRS.
Ethics. All user studies are completed by anonymous participation of volunteers to avoid sensitive data disclosure. Approval of all ethical and experimental procedures and protocols was granted by Science and Technology Ethics Committee of Zhejiang University.
@@
    ●, ◐, ○, from high to low, represent different matching degrees; ”-” means no requirement.
@@ 2 RELATED WORK
2.1 Face Morphing Attacks.
Face morphing attack [15] involves creating a digitally composite or ”morphed” face image. The ”morphed” face combines the facial features of two or more individuals of different identities, allowing the attackers to bypass authentication processes after enrolling the generated facial image into the database. Morphing attacks can be conducted in either the image or feature space. The former generates a forged image by interpolating the facial images of different identities together [16, 17]. The latter aims to reconstruct an average face to make it close to the given identities in feature space, using a generative adversarial network [18, 16] for generating morphed face. However, the performance of these attacks is limited by the number of identities that need to be fused. For example, Wu et al. [4] introduced UniID, a method to inject a specialized disguised face during the enrollment phase, enabling multiple attackers to bypass the FRS. Since UniID necessitates the facial images of the attackers to generate the patch to disguise the enrolled face, its performance diminishes nearly to zero as the number of attackers increases to 10. Li et al. [12] presented an attack that directly injected the backdoor by modulating LED in a specialized waveform during the enrollment phase, raising the assumption for adversaries’ capabilities. Besides, their method requires accessibility to facial images of the attackers, which can thus be viewed as a morphing attack.
2.2 Adversarial Attacks against FRS.
@@
where $f_{\theta}(x_{t})$ is the feature vectors of targeted identity for optimization and $l(\cdot)$ is the overall loss function to satisfy attack constraints. This goal implies that for each pair of attacker and victim $(x_{a},x_{t})$, the attacker must generate a distinctive patch $p$, entailing a significant attack cost [6, 19]. The concept of a universal adversarial attack has been proposed to enable adversarial perturbations to be universally applicable across attackers with different identities. Amada et al. [20] utilized small, pixel-level perturbations that, when added to facial images, can deceive the FRS into recognizing a face as belonging to several distinct identities simultaneously. Yang et al. [21] developed ”universal adversarial patch” that can reliably fool face detectors into failing to detect the faces. However, these adversarial attacks are only launched in the authentication stage, and their generalization and universality are limited by the difference between the surrogate and target model.
2.3 Backdoor Attacks against FRS.
Backdoor attacks against FRS entail manipulating the model’s training data to introduce a concealed vulnerability that attackers can exploit in the authentication stage. This attack aims to compromise the system’s integrity and enable unauthorized access to bypass facial recognition authentication. Guo et al. [10] introduced a hidden backdoor called ”master key” into the feature extractor during its training phase and activated the backdoor using a specific trigger. These attacks [10, 22, 7] focus on compromising the FRS by poisoning the training data of the feature extractors. Nevertheless, such an endeavor is frequently unrealistic, given that access to the training data is often restricted.
@@ 2.4 Remarks
Systematically, we summarize our attack and list the differences of previous techniques from several dimensions that matter a lot to launch attacks in real-world scenarios.
Feasibility. Feasibility denotes whether the attacks can be launched easily in real-world scenarios. Backdoor attacks [10, 22, 7] requiring the manipulation of the training data is far from feasible since FRS are often unknown for attackers. Adversarial attacks [23, 24] need extra computations for each new attacker, assuming high computational resources for the adversary. As for face morphing [16, 17] and FIBA, the only step that needs computation is to generate the ”average or triggered face” and enroll it in the database.
Universality. Universality means whether the attacks are universally applicable to any attackers when attackers’ facial images are out of reach. This characteristic measures the severity of the threat since high universality brings large-scale and rapid spread. Face morphing attacks [16, 17, 16] that need attackers’ facial images to generate morphed faces have little universality. Adversarial attacks [25, 26, 27] have minor universality because of their transferability. Backdoor attacks [10, 22, 7] and FIBA do not specify attackers and are universally applicable.
Generalization. Generalization refers to whether the attacks are effective to different target FRS under the most practical (black-box) settings, with no accessibility to models of FRS. Since both morphing and adversarial attacks [25, 17] rely largely on the surrogate model, their generalization is limited. For the black-box FRS, backdoor attacks [10, 22, 7] have no authority to modify the training data of the pretrained recognition models as well. Note that under the black-box setting, FIBA hardly depends on models, with high transferability which is validated in the evaluation.
Stealthiness. Stealthiness measures whether the attacks will be detected since advanced FRS have deployed with FLV and other anomaly detection methods. Face morphing attacks [18, zhang2021mipgan, 16] often optimize in the digital space to generate a morphed face without any stealthiness to evade FLV. However, the other three attacks [28, 29, 30] can be launched with high flexibility to avoid potential detection.
@@ 3.1 Face Recognition Systems
The initial face detection stage aims to localize and isolate regions within the input image that potentially contain a human face and the next stage aims to localize key facial landmarks. With facial landmarks established, the system can normalize the detected face by applying geometric transformations like scaling, rotation, and alignment. Finally, having obtained these normalized face representations: (a) during enrollment phase, the extracted representations will be stored in the feature database for retrieval; (b) during authentication phase, the face matching stage compares the extracted representations with features in the database.
@@
$A=\{id|\mathrm{cos}(f_{\theta}(x),y_{id})>\delta_{match},y_{id}\in\mathbb{D}\}$ (2)
where $x$ is a normalized face image, $\mathrm{cos}(\cdot)$ denotes cosine similarity and $\delta_{match}$ is a threshold for face matching. If the result $A=\phi$, the identity does not exist in $\mathbb{D}$, and vice versa. Therefore, $f_{\theta}$ is the key component for FRS to distinguish various identities. To successfully spoof the FRS, it is crucial to deceive $f_{\theta}$ first.
@@ 3.2 Inherent Vulnerability of FRS
With the preliminaries of FRS, we are going to investigate the vulnerabilities in real-world settings. Starting from our own experience (often misidentified as others), we notice that FRS is susceptible not only to deliberate malicious attacks but also to various environment variables, significantly affecting user experience. To obtain a more comprehensive understanding of this phenomenon, a user survey is conducted to collect instances of mistaken identity recognition or failure to recognize users during FRS utilization. 80 volunteers (details presented in Appendix B.4) are recruited, and each participant is administered a questionnaire. It should be noted that 73 participants have prior experience using FRS. The user survey results are depicted in Figure 3. For those who have used FRS before, a staggering 47.95% of participants reported instances of being misidentified as someone else, and an overwhelming 80.82% experienced failures to be recognized by the system, highlighting that there are inherent flaws within FRS, presenting a significant obstacle to widespread adoption and seamless user experience. These naturally occurring evasion and impersonation incidents suggest that FRS is not as reliable as previously assumed, even without deliberate attacks from adversaries. These findings act as a sobering wake-up call, prompting an in-depth investigation into these issues’ root causes and ramifications. A pressing question demand our attention:
@@
Figure 3: Results of user study. Question I: Have you ever been mistaken for someone else when using FRS? Question II: Have you ever been unrecognized when using FRS?

### v3

@@ Abstract
Face Recognition Systems (FRS) have increasingly integrated into critical applications, including surveillance and user authentication, highlighting their pivotal role in modern security systems. Recent studies have revealed vulnerabilities in FRS to adversarial (e.g., adversarial patch attacks) and backdoor attacks (e.g., training data poisoning), raising significant concerns about their reliability and trustworthiness. However, previous studies have been based on the overclaimed adversary’s capabilities, with limited feasibility in practice.
Correspondingly, in this paper, we delve into the inherent vulnerabilities in FRS through user studies and preliminary explorations. By exploiting these vulnerabilities, we identify a novel attack, Facial Identity Backdoor Attack dubbed FIBA, which unveils a potentially more devastating threat against FRS: an enrollment-stage backdoor attack. FIBA enable broad-scale disruption by allowing any attacker donning a specific trigger to bypass these systems. This implies that after a single, poisoned example is enrolled into the database, the corresponding trigger becomes a universal key for any attackers to spoof the FRS. This strategy essentially challenges the conventional attacks by initiating at the enrollment stage, dramatically transforming the threat landscape by poisoning the feature database rather than the training data. We conduct extensive evaluations encompassing digital and physical experiments across 6 face recognition models, 5 commercial face authentication APIs, and 3 IoT devices to demonstrate FIBA’s priorities over traditional attacks. Across these scenarios, FIBA achieves an attack success rate peaking at 100%, underscoring its potential for widespread exploitation. Additionally, the impacts of various physical factors are also discussed and evaluated. Experimental findings together with our suggestions are reported to affected vendors. We conclude by proposing initial mitigation strategies for service vendors and offer a framework for preliminary defense against FIBA with evaluation results.
@@ 1 Introduction
Face recognition systems (FRS) are computer applications that identify or verify a person’s identity from facial images or video frames by analyzing the facial feature patterns [1, 2]. These systems have become increasingly important across multiple domains owing to their efficient and non-invasive identification and authentication capabilities. For instance, between 2017 and 2019, 64 countries adopted artificial intelligence surveillance through FRS [3] and 7 in 10 governments use FRS extensively, among which 20% use it on some buses, 30% use it on trains or subways, and about 60% use it in some airports [3]. Recent researches [4, 5, 6] have demonstrated that these security-sensitive systems are vulnerable to adversarial patch attacks achieved through elaborate patches or other disguises during the authentication. Traditional backdoor attacks [7, 8] can backdoor FRS by poisoning the training data of the face recognition models in advance. However, these attacks are founded on assumptions that attackers are capable of generating different patches with intensive computation resources for each source-target pair or manipulating the training data. These assumptions result in large gaps of adversary’s capabilities between research and practice [9, 10, 11], substantially increasing the cost for adversaries in real-world scenarios.
@@
Figure 1: Illustration of FIBA. An insider wearing a trigger registers his/her face into the database. During authentication, attackers with the same trigger can bypass FRS.
@@
From the perspective of attackers, we wonder that “If it is possible to spoof a targeted FRS with practical adversary’s capabilities”. Under this practical scenario, details of the FRS (e.g., training data of the feature extractor) are out of reach, launching resource-intensive attacks is impractical and so on. Besides, considering the adversary’s goal, attack performance (e.g., applicable to any attackers, stealthy to evade detection, generalizable to different FRS and robust in physical world) under complex scenarios is also important. As such, executing such an attack is far from trivial. To achieve the mentioned attack goals and further gain more insights into the vulnerabilities inherent in FRS, we first conducted a user study, with 80 volunteers involved, focused on naturally occurring impersonation and evasion attacks in real-world scenarios. The results indicate that, despite claims of accuracy exceeding 99% by numerous facial recognition service vendors, 47.95% and 80.82% of FRS users have experienced misidentification or outright rejection by these services, respectively. After that, to figure it out, we conduct preliminary experiments and reveal the inherent flaws of FRS, that face recognition models rely on a small subset of features to distinguish two facial images of different identities. By exploiting these flaws, we then introduce an enrollment-stage Facial Identity Backdoor Attack (FIBA)–a novel attack vector that allows attackers to bypass FRS by enrolling an insider’s face disguised with a trigger into the system, without the requirement for training data accessibility. Attackers possessing the trigger can subsequently fool the FRS with ease. Note that an insider denotes the one who has the ability to register his/her face into the database, a common privilege in FRS [12, 13] (see in Section 4). The overall attack process is shown in Figure 1. The underlying mechanism is that distinguished facial features can be substituted with our meticulously crafted features to serve as a backdoor. Faces disguised with the same trigger will represent similar facial features.
@@
Even with the help of a compromised insider to register a facial image into the database of FRS, launching FIBA in real-world scenarios is still challenging since the adversaries have minor knowledge about FRS, e.g., inaccessibility of the pre-trained models of commercial APIs or IoT devices. In such a black-box setting, ensuring the generalizability of FIBA across different FRS presents significant challenges. Furthermore, bypassing FRS also entails evading facial liveness verification (FLV), a task that is likewise nontrivial. Additionally, different from previous attacks [4, 14], the attacker’s images remain inaccessible, necessitating its universality towards various identities. The robustness of FIBA is also vital for resisting the effects of digital-to-physical distortion that may degrade the attack performance. Consequently, it is imperative to address the following challenges:
@@
To address challenge 1), a data-dependent backdoor feature optimization algorithm is introduced to eliminate the parameter disparities and other differences among the facial feature extractors. This algorithm leverages the common vulnerabilities within FRS to optimize the generation of the patch (trigger) with key facial features, effectively mitigating disparities caused by model parameters and structures. To safeguard the attack’s stealthiness and effectiveness in challenge 2), exploratory experiments are conducted to pinpoint the locations of key facial features in advance, ensuring the minimization of the facial alteration area while maximizing attack performance. To fulfill the universality in challenge 3), attacker-agnostic patch optimization flow is used by applying the same patch for both attackers and insiders. Finally, to address challenge 4), external dataset augmentation and physical transformation techniques are employed to guarantee the robustness of generated triggers to digital-to-physical distortion. With the above challenges addressed, FIBA can breach practical applications easily. The following is an example attack scenario: for a university deployed with FRS, a student can enroll his/her disguised face and distribute the trigger for profit (note that enrolling or updating facial images are common in practice, e.g., HikVision [13] and Invixium [12]), allowing anyone with the same trigger to bypass the FRS. Since the trigger is applicable to anyone, this threat can be amplified rapidly, as the trigger is spread to more people. FIBA can also introduce similar threats in companies, housing estates, and other departments equipped with FRS (i.e., imagine that everyone could bypass the FRS of a security-critical department).
@@
  • Based on our analysis, we have reported our experimental findings to affected vendors together with our suggestions. A potential defense strategy is proposed to mitigate the threat of FIBA, and a preliminary evaluation is conducted to shed light on building more reliable FRS.
Ethics. All user studies are completed by anonymous participation of volunteers to avoid sensitive data disclosure. Approval of all ethical and experimental procedures and protocols was granted by Science and Technology Ethics Committee of XXX University.
@@
    ●, ◐, ○, from high to low, represent different matching degrees; “-” means no requirement.
@@ 2 RELATED WORK
2.1 Face Morphing Attacks
Face morphing attack [15] involves creating a digitally composite or “morphed” face image. The “morphed” face combines the facial features of two or more individuals of different identities, allowing the attackers to bypass authentication processes after enrolling the generated facial image into the database. Morphing attacks can be conducted in either the image or feature space. The former generates a forged image by interpolating the facial images of different identities together [10, 16]. The latter aims to reconstruct an average face to make it close to the given identities in feature space, using a generative adversarial network [9, 10] for generating morphed face. However, the performance of these attacks is limited by the number of identities that need to be fused. For example, Wu et al. [4] introduced UniID, a method to inject a specialized disguised face during the enrollment phase, enabling multiple attackers to bypass the FRS. Since UniID necessitates the facial images of the attackers to generate the patch to disguise the enrolled face, its performance diminishes nearly to zero as the number of attackers increases to 10. Li et al. [14] presented an attack that directly injected the backdoor by modulating LED in a specialized waveform during the enrollment phase, raising the assumption for adversaries’ capabilities. Besides, their method requires accessibility to facial images of the attackers, which can thus be viewed as a morphing attack.
2.2 Adversarial Attacks against FRS
@@
where $f_{\theta}(x_{t})$ is the feature vectors of targeted identity for optimization and $l(\cdot)$ is the overall loss function to satisfy attack constraints. This goal implies that for each pair of attacker and victim $(x_{a},x_{t})$, the attacker must generate a distinctive patch $p$, entailing a significant attack cost [6, 17]. The concept of a universal adversarial attack has been proposed to enable adversarial perturbations to be universally applicable across attackers with different identities. Amada et al. [18] utilized small, pixel-level perturbations that, when added to facial images, can deceive the FRS into recognizing a face as belonging to several distinct identities simultaneously. Yang et al. [19] developed a “universal adversarial patch” that can reliably fool face detectors into failing to detect the faces. However, these adversarial attacks are only launched in the authentication stage, and their generalization and universality are limited by the difference between the surrogate and target model.
2.3 Backdoor Attacks against FRS
Backdoor attacks against FRS entail manipulating the model’s training data to introduce a concealed vulnerability that attackers can exploit in the authentication stage. This attack aims to compromise the system’s integrity and enable unauthorized access to bypass facial recognition authentication. Guo et al. [20] introduced a hidden backdoor called “master key” into the feature extractor during its training phase and activated the backdoor using a specific trigger. These attacks [20, 21, 7] focus on compromising the FRS by poisoning the training data of the feature extractors. Nevertheless, such an endeavor is frequently unrealistic, given that access to the training data is often restricted.
@@ 2.4 Remarks
Systematically, we summarize our attack and list the differences between previous techniques from several dimensions that matter a lot in real-world scenarios.
Feasibility. Feasibility denotes whether the attacks can be launched easily in real-world scenarios. Backdoor attacks [20, 21, 7] requiring the manipulation of the training data is far from feasible since FRS are often unknown to attackers. Adversarial attacks [22, 23] need extra computations for each new attacker, assuming high computational resources for the adversary. As for face morphing [10, 16] and FIBA, the only step that needs computation is to generate the “average or triggered face” and enroll it in the database.
Universality. Universality means whether the attacks are universally applicable to any attackers when attackers’ facial images are out of reach. This characteristic measures the severity of the threat since high universality brings large-scale and rapid spread. Face morphing attacks [10, 16] that need attackers’ facial images to generate morphed faces have little universality. Adversarial attacks [24, 25, 26] have minor universality because of their transferability. Backdoor attacks [20, 21, 7] and FIBA do not specify attackers and are universally applicable.
Generalization. Generalization refers to whether the attacks are effective to different target FRS under the most practical (black-box) settings, with no accessibility to models of FRS. Since both morphing and adversarial attacks [24, 16] rely heavily on the surrogate model, their generalization is limited. For the black-box FRS, backdoor attacks [20, 21, 7] have no authority to modify the training data of the pretrained recognition models as well. Note that under the black-box setting, FIBA hardly depends on models, with high transferability which is validated in the evaluation.
Stealthiness. Stealthiness measures whether the attacks will be detected since advanced FRS have deployed with FLV and other anomaly detection methods. Face morphing attacks [9, 10] often optimize in the digital space to generate a morphed face without any stealthiness to evade FLV. However, the other three attacks [27, 28, 29] can be launched with high flexibility to avoid potential detection.
@@ 3.1 Face Recognition Systems
The initial face detection stage aims to localize and isolate regions within the input image that potentially contain a human face and the next stage aims to localize key facial landmarks. With facial landmarks established, the system can normalize the detected face by applying geometric transformations like scaling, rotation, and alignment. Finally, having obtained these normalized face representations: (a) during the enrollment phase, the extracted representations will be stored in the feature database for retrieval; (b) during the authentication phase, the face matching stage compares the extracted representations with features in the database.
@@
$A=\{\mathrm{id}|\mathrm{cos}(f_{\theta}(x),y_{\mathrm{id}})>\mathrm{\delta_{match}},y_{id}\in\mathbb{D}\}$ (2)
where $x$ is a normalized face image, $\mathrm{cos}(\cdot)$ denotes cosine similarity and $\mathrm{\delta_{match}}$ is a threshold for face matching. If the result $A=\phi$, the identity does not exist in $\mathbb{D}$, and vice versa. Therefore, $f_{\theta}$ is the key component for FRS to distinguish various identities. To successfully spoof the FRS, it is crucial to deceive $f_{\theta}$ first.
@@ 3.2 Inherent Vulnerability of FRS
With the preliminaries of FRS, we are going to investigate the vulnerabilities in real-world settings. Starting from our own experience (often misidentified as others), we notice that FRS is susceptible not only to deliberate malicious attacks but also to various environment variables, significantly affecting user experience. To obtain a more comprehensive understanding of this phenomenon, a user survey is conducted to collect instances of mistaken identity recognition or failure to recognize users during FRS utilization. 80 volunteers (details presented in Appendix 5) are recruited, and each participant is administered a questionnaire. It should be noted that 73 participants have prior experience using FRS. The user survey results are depicted in Figure 3. For those who have used FRS before, a staggering 47.95% of participants reported instances of being misidentified as someone else, and an overwhelming 80.82% experienced failures to be recognized by the system, highlighting that there are inherent flaws within FRS, presenting a significant obstacle to widespread adoption and seamless user experience. These naturally occurring evasion and impersonation incidents suggest that FRS is not as reliable as previously assumed, even without deliberate attacks from adversaries. These findings act as a sobering wake-up call, prompting an in-depth investigation into these issues’ root causes and ramifications. A pressing question demands our attention:
@@
Figure 3: Results of user study. Question 1: Have you ever been mistaken for someone else when using FRS? Question 2: Have you ever been unrecognized when using FRS?